Given 1.25 USD/GBP, 1.10 USD/EUR, and .9000 GBP/EUR, arbitrage to earn USD. Assume you've 1m USD line of credit with 0% interest rate.

  • A. $22,727.27
  • B. $27,272.73
  • C. $18,272.27
  • D. $18,727.73
Answer #1
  • We have 10,00,000 USD
  • Convert to EUR ie

1.1 USD = 1 EUR

1 USD = 1/1.1 EUR

10,00,000 USD = (1/1.1)*10,00,000 EUR

= 909090.91 EUR

  • Convert EUR to GBP ie

.9 GBP = 1 EUR

909090.91 EUR = .9*909090.91 GBP

= 818181.82 GBP

  • Convert GBP to USD ie

1.25 USD = 1 GBP

818181.82 GBP = 1.25* 818181.82 USD

= 1022727.275 USD

  • Arbitrage prift = 1022727.275-1000000 = ​​​​​​​$22727.27